Firms HL and LL are identical except for their financial leverage ratios and the interest rates they pay on debt. Each has $10 million in invested capital, has $1.5 million of EBIT, and is in the 25% federal-plus-state tax bracket. Firm HL, however, has a debt-to-capital ratio of 55% and pays 11% interest on its debt, whereas LL has a 25% debt-to-capital ratio and pays only 10% interest on its debt. Neither firm uses preferred stock in its capital structure.

Calculate the return on invested capital (ROIC) for each firm. Round your answers to two decimal places

. ROIC for firm LL: %

ROIC for firm HL: %

Calculate the rate of return on equity (ROE) for each firm. Round your answers to two decimal places.

ROE for firm LL: %

ROE for firm HL: %

Observing that HL has a higher ROE, LL's treasurer is thinking of raising the debt-to-capital ratio from 25% to 60% even though that would increase LL's interest rate on all debt to 15%. Calculate the new ROE for LL. Round your answer to two decimal places. %
Answer #1

a.ROIC = EBIT(1-T)/Capital

LL = 1.5 million(1-25%)/10 million = 11.25%

HL = 1.5 million(1-25%)/10 million = 11.25%

b.Return on Equity = Income for Equity/Equity Funds

LL = (1.5 million – 0.25)(1-25%)/7.5

= 12.5%

HL = (1.5 – 0.605)(1-25%)/4.5

= 14.92%

c.ROE = (1.5 – 0.9)(1-25%)/4

= 11.25%